# Reward Engineering for Spatial Epidemic Simulations: A Reinforcement Learning Platform for Individual Behavioral Learning

### v1

@@ 5 Results
Performance of different algorithms in ContagionRL environment. To evaluate the utility of the ContagionRL for obtaining effective learned policies via reinforcement learning (RL), we trained and assessed three distinct RL algorithms: Soft Actor-Critic (SAC) (schulman2017proximal), Proximal Policy Optimization (PPO) (schulman2017proximal), and Advantage Actor-Critic (A2C) (a3c_paper_mnih). Their performance was benchmarked against baselines mentioned in section 4 (configuration details are in Appendix C.2 and Appendix C.4).
@@
An important observation, across all tested reward functions, is the consistent emergence of high NPI adherence in the learned policies. This suggests that the benefit of maximizing adherence is a relatively identifiable signal within the ContagionRL environment, discoverable through learning irrespective of the specific reward structure guiding navigation. The primary challenge lies in learning effective spatial navigation strategies, a task for which the choice of reward function is more important. The data presented in Figures 3 and 10, and substantiated by the statistical comparisons in Table 8, unequivocally demonstrate the influence of reward function design on the learned policy to achieve the desired infection avoidance behavior.
@@
Comparison of Partial Observability (POMDP) and Full Observability. To understand the learned behavior under limited visibility, we formulate the epidemic control problem as a partially observable Markov decision process (POMDP). In real epidemics, actors have incomplete information about the environment state and infection status of individuals (gersovitz2004economical; geoffard1996rational; farboodi2021internal). We implement this POMDP formulation using a visibility radius constraint, where individuals within radius $r$ of RL agent’s position are passed as observations and individuals beyond this radius are unobservable. This natural information asymmetry captures a realistic scenario where agents have limited sensing or information networks that decay with distance. Figure 5 demonstrates trained partial visibility models (r=10, r=15, r=20) paradoxically outperform the baselines and the full visibility trained model, Trained Full, in average reward, episode length and infections per timestep. This counter-intuitive result suggests that complete information can be detrimental to learning, potentially due to observation noise and the increased dimensionality in high-dimensional observation spaces. Limited observability forces the agent to develop more robust strategies and reduces the complexity of the decision space. The trained (Full) model’s performance is on par with the greedy heuristic. Notably, a performance differences between r=10, r=15 and r=20 conditions are not statistically significant according to Table 10, indicating a performance plateau beyond a certain visibility threshold.
@@
Social and Economic Constraint in Human Movement. Mobility is a key consideration in epidemiological models (Fakir2021) and literature review shows that mobility patterns remain unchanged in a pandemic for essential or low wage workers (Gallagher2021). Inspired by this, we conducted an experiment to evaluate the performance of the agent under these socioeconomic constraints in mobility. Figure 6 shows that the trained PPO agent outperforms stationary, random and greedy baselines when trained on a workplace/home movement cycle. We notice that the stationary and random baselines declined in this cyclic movement, but the trained and greedy baseline improve considerably. Furthermore, the trained RL agent has a higher average reward and higher mean episode length, with an extremely low rate of infections per timestep (Table 2). In the random environment, the agent struggles with long predictions and has a short horizon on its reward, as it performs worse than the greedy baseline. We attribute its inability to beat the local optimization of the greedy algorithm to the noisy observations and movements of the random environment. By incorporating the spatio-temporal regularities of a workplace/home mobility cycle, a trained agent can formulate more effective long-term strategies. The resulting expansion of its planning horizon is empirically validated by a substantial improvement in cumulative average reward over an agent exposed to a random mobility pattern. We anticipated that trained agents operating under cyclic movement patterns would exhibit lower standard deviation across performance metrics compared to those under random movement. However, our results contradicted this expectation. Furthermore, while infections per timestep increased for all agent types, the trained agent demonstrated no such increase, maintaining stable infection rates throughout the evaluation period.
Figure 6: Performance comparison of RL agents across different human movement patterns in epidemic control. The figure shows three metrics across two movement patterns: Random Movement (continuous random walk) and Workplace/Home Cycle (structured movement between workplace and home locations). Four agent types are compared: Stationary, Random, Greedy, and Trained (PPO agent). Results are averaged across 3 random seeds with 100 evaluat ion episodes per seed. Error bars show 95% bootstrap confidence intervals from per-seed means. Left: Average cumulative reward per episode. Center: Mean episode length (survival time). Right: Infection rate. Trained agent significantly outperform all baselines in the Workplace/Home Cycle environment but show mixed results in Random Movement. Statistical results can be found in Table 2
@@
Workplace/Home
Cycle
Trained vs Stationary Trained 89,067.0 $p<0.001$ ***
@@ A.3 Comparison of the epidemiological models
SIR (kermack1927contribution) S, I, R No No No None No
SIRS (kermack1927contribution) S, I, R Yes No No None No
SEIR (kermack1927contribution; wilson1945law) S, E, I, R No Yes No None No
SIRS+D (wilson1945law) S, I, R, D Yes No Yes None No
ABM (grimm_2013_abm) Custom per agent Yes Optional Yes Prescribed rules Partial
@@ C.3.5 Potential Field Reward
C.4 Algorithmic Comparison Hyperparameters
@@
C.5 Baseline Implementation Details
@@
C.5.1 Stationary Policy Implementation
@@
C.5.2 Random Policy Implementation
@@
C.5.3 Greedy Distance Maximizer Policy Implementation
@@ Appendix D Results Continued
Trained vs
Baselines
Full vs Stationary Full 79,732.0 $p<0.001$ ***
@@
Trained vs
Trained
Full vs r=10 r=10 35,009.0 $p<0.001$ ***
@@
Appendix E Compute Resources
@@
Appendix F Additional Results and Experiments

### v2

@@ 5 Results
Performance of different algorithms in ContagionRL environment. To evaluate the utility of the ContagionRL for obtaining effective learned policies via reinforcement learning (RL), we trained and assessed three distinct RL algorithms: Soft Actor-Critic (SAC) (Schulman et al., 2017), Proximal Policy Optimization (PPO) (Schulman et al., 2017), and Advantage Actor-Critic (A2C) (Mnih et al., 2016). Their performance was benchmarked against baselines mentioned in section 4 (configuration details are in Appendix C.2 and Appendix C.5).
@@
An important observation, across all tested reward functions, is the consistent emergence of high NPI adherence in the learned policies. This suggests that the benefit of maximizing adherence is a relatively identifiable signal within the ContagionRL environment, discoverable through learning irrespective of the specific reward structure guiding navigation. A detailed analysis of learned adherence behavior across all reward functions and ablations is provided (Appendix E). The primary challenge lies in learning effective spatial navigation strategies, a task for which the choice of reward function is more important. The data presented in Figures 3 and 10, and substantiated by the statistical comparisons in Table 10, unequivocally demonstrate the influence of reward function design on the learned policy to achieve the desired infection avoidance behavior.
@@
[…]
vs Random vs Static
Policy New Inf./step $\Delta$% Sig. $\Delta$% Sig.
Random (baseline) 1.133
Static (baseline) 1.121
Potential Field 0.890 -21.4% ** -20.6% **
Const.+Red. Inf. 0.925 -18.3% ** -17.5% **
Constant 0.963 -15.0% n.s. -14.2% *
Max Near. Dist. 0.972 -14.2% n.s. -13.3% n.s.
Reduce Infection 1.007 -11.1% n.s. -10.2% n.s.
@@
Comparison of Partial Observability (POMDP) and Full Observability. To understand the learned behavior under limited visibility, we formulate the epidemic control problem as a partially observable Markov decision process (POMDP). In real epidemics, actors have incomplete information about the environment state and infection status of individuals (Gersovitz and Hammer, 2004; Geoffard and Philipson, 1996; Farboodi et al., 2021). We implement this POMDP formulation using a visibility radius constraint, where individuals within radius $r$ of RL agent’s position are passed as observations and individuals beyond this radius are unobservable. This natural information asymmetry captures a realistic scenario where agents have limited sensing or information networks that decay with distance. Figure 5 demonstrates trained partial visibility models (r=10, r=15, r=20) paradoxically outperform the baselines and the full visibility trained model, Trained Full, in average reward, episode length and infections per timestep. This counter-intuitive result suggests that complete information can be detrimental to learning, potentially due to observation noise and the increased dimensionality in high-dimensional observation spaces (see Appendix F for feature importance analysis). Limited observability forces the agent to develop more robust strategies and reduces the complexity of the decision space. The trained (Full) model’s performance is on par with the greedy heuristic. Notably, a performance differences between r=10, r=15 and r=20 conditions are not statistically significant according to Table 12, indicating a performance plateau beyond a certain visibility threshold. We note that the potential field reward is computed from global state even under POMDP conditions (Section 3.3), functioning as a privileged training signal. The agent’s improved performance under limited visibility therefore reflects genuine policy robustness rather than information leakage through observations.
@@
Social and Economic Constraint in Human Movement. Mobility is a key consideration in epidemiological models (Fakir and Bharati, 2021) and literature review shows that mobility patterns remain unchanged in a pandemic for essential or low wage workers (Gallagher et al., 2021). Inspired by this, we conducted an experiment to evaluate the performance of the agent under these socioeconomic constraints in mobility. Figure 6 shows that the trained PPO agent outperforms stationary, random and greedy baselines when trained on a workplace/home movement cycle. We notice that the stationary and random baselines declined in this cyclic movement, but the trained and greedy baseline improve considerably. Furthermore, the trained RL agent has a higher average reward and higher mean episode length, with an extremely low rate of infections per timestep (Table 3). In the random environment, the agent struggles with long predictions and has a short horizon on its reward, as it performs worse than the greedy baseline. We attribute its inability to beat the local optimization of the greedy algorithm to the noisy observations and movements of the random environment. By incorporating the spatio-temporal regularities of a workplace/home mobility cycle, a trained agent can formulate more effective long-term strategies. The resulting expansion of its planning horizon is empirically validated by a substantial improvement in cumulative average reward over an agent exposed to a random mobility pattern. We anticipated that trained agents operating under cyclic movement patterns would exhibit lower standard deviation across performance metrics compared to those under random movement. However, our results contradicted this expectation. Furthermore, while infections per timestep increased for all agent types, the trained agent demonstrated no such increase, maintaining stable infection rates throughout the evaluation period.
Figure 6: Performance comparison of RL agents across different human movement patterns in epidemic control. The figure shows three metrics across two movement patterns: Random Movement (continuous random walk) and Workplace/Home Cycle (structured movement between workplace and home locations). Four agent types are compared: Stationary, Random, Greedy, and Trained (PPO agent). Results are averaged across 3 random seeds with 100 evaluation episodes per seed. Error bars show 95% bootstrap confidence intervals from per-seed means. Left: Average cumulative reward per episode. Center: Mean episode length (survival time). Right: Infection rate. Trained agents significantly outperform all baselines in the Workplace/Home Cycle environment but show mixed results in Random Movement. Statistical results can be found in Table 3
@@
Workplace/Home Cycle Trained vs Stationary Trained 89,067.0 $p<0.001$ ***
@@ 6 Conclusion
Broader Impact Statement. ContagionRL provides a platform for researchers to study how reward design influences learned behaviour in an epidemic setting. In a broad sense, experiments performed using platforms of this type can help us understand how potential public health intervention strategies, leading to perceived rewards or avoiding penalties, might shape protective behaviours. Nonetheless, we emphasize that human behaviour around disease avoidance is complex and heterogeneous, and that the relationship between simulation-encoded rewards and real-world perceived rewards is also challenging to understand. Simulation findings will complement rather than replace epidemiological expertise and empirical validation for the foreseeable future.
@@ A.3 Comparison of the epidemiological models
SIR (Kermack and McKendrick, 1927) S, I, R No No No None No
SIRS (Kermack and McKendrick, 1927) S, I, R Yes No No None No
SEIR (Kermack and McKendrick, 1927; Wilson and Worcester, 1945) S, E, I, R No Yes No None No
SIRS+D (Wilson and Worcester, 1945) S, I, R, D Yes No Yes None No
ABM (Grimm and Railsback, 2013) Custom per agent Yes Optional Yes Prescribed rules Partial
@@ C.3.5 Potential Field Reward
C.4 Reward Function Information Scope
The summation in $\mathbf{F}$ iterates over all humans in the simulation irrespective of the agent’s visibility radius parameter. Consequently, under POMDP configurations where the agent’s observation is restricted to a radius $r$, the force vector $\mathbf{F}$ still incorporates contributions from humans beyond $r$. The observation function zeroes out features for humans outside the visibility radius, but the reward computation does not apply this filter. This asymmetry means the directional reward acts as an oracle-like gradient signal during training, informing the agent whether its action aligned with the globally optimal escape direction even when the agent could not observe all contributing humans. This is consistent with asymmetric actor-critic approaches where training-time information may exceed test-time observations (Pinto et al., 2017).
Table 6: Information scope of each reward function and its components. Local terms depend only on the agent’s own state or action. Global terms are computed from non-agent humans regardless of the agent’s visibility radius (under the default MDP configuration with visibility_radius = 1-1).
Reward Function Component Scope Depends On
Constant Alive bonus Local Agent state ($S_{a}$)
Reduce Infection Alive check Local Agent state ($S_{a}$)
$(1-P_{\text{inf}})^{2}$ Global Agent pos., adherence, all infected positions
Const. + Red. Inf. Alive bonus Local Agent state ($S_{a}$)
Alive check Local Agent state ($S_{a}$)
$(1-P_{\text{inf}})^{2}$ Global Agent pos., adherence, all infected positions
Max Near. Dist. Alive check Local Agent state ($S_{a}$)
$d_{\min}/D_{\beta}$ Global Agent pos., all S & I positions
Potential Field Health ($r_{\text{health}}$) Local Agent state ($S_{a}$)
Adherence ($r_{\text{adherence}}$) Local Agent action ($\alpha$)
Direction ($r_{\text{dir}}$) Global Agent action, all S & I positions
Magnitude ($r_{\text{mag}}$) Global Agent action, all S & I positions
C.5 Algorithmic Comparison Hyperparameters
@@
C.6 Baseline Implementation Details
@@
C.6.1 Stationary Policy Implementation
@@
C.6.2 Random Policy Implementation
@@
C.6.3 Greedy Distance Maximizer Policy Implementation
@@ Appendix D Results Continued
Trained vs Baselines Full vs Stationary Full 79,732.0 $p<0.001$ ***
@@
Trained vs Trained Full vs r=10 r=10 35,009.0 $p<0.001$ ***
@@
Appendix E Learned Adherence Behavior Analysis
To investigate adherence usage of trained models across different reward formulations, we conducted an analysis of the learned adherence behaviors. This analysis examines if agents can independently discover the value of non-pharmaceutical intervention (NPI) compliance or if they require explicit reward tuning for discovering the importance of adherence.
Methodology. We measured the mean and standard deviation of adherence actions from trained policies across all reward functions and potential field ablation variants. Each policy was evaluated using the same experimental setup as the main results. Adherence values were recorded at each timestep and aggregated across episodes.
Results. Table 13 shows the learned adherence usage across reward functions. Agents trained with simple reward functions like Constant, Reduce Infection, and Constant + Reduce Infection achieve maximal adherence ($\alpha=1.0$) with minimal variance. This indicates that even without explicit adherence incentives, temporal credit assignment enables agents to discover the survival benefit of full NPI compliance through the sparse reward signal. The Max Distance reward function yields slightly lower mean adherence with higher variance, suggesting that spatial optimization partially competes with adherence learning in this formulation.
The Full Potential Field reward achieves maximal adherence, consistent with its explicit adherence component. However, when the adherence reward is removed from potential field reward function (PF: No Adherence), mean adherence drops to 0.41 with high variance of 0.47. This indicates essentially random adherence behavior. All other potential field ablations (No Direction, No Health, No Magnitude, No Movement and No Susceptible Repulsion) maintain maximal adherence despite their respective component removals.
These results reveal an asymmetry in adherence learnability based on reward complexity. Simple reward functions allow agents to implicitly learn maximal adherence through the survival signal alone. However, in complex multi-component objectives like the potential field reward, agents fail to discover adherence value without explicit incentivization. The diverse reward signals in the potential field formulation appear to dilute the implicit survival gradient that simpler rewards preserve.
The ContagionRL platform includes a configurable adherence penalty parameter (Table 5) to enable the future investigation of strategic trade-offs where agents must balance survival benefits against adherence costs.
Table 13: Learned adherence behavior across reward functions and potential field ablations. Mean and standard deviation computed over 300 evaluation episodes (3 seeds $\times$ 100 episodes). PF denotes Potential Field.
Reward Function Mean Adherence Std Adherence
Constant 1.000 0.000
Reduce Infection 1.000 0.000
Constant + Reduce Inf 1.000 0.000
Max Distance 0.872 0.312
Potential Field 1.000 0.001
PF: No Adherence 0.406 0.472
PF: No Direction 1.000 0.000
PF: No Health 1.000 0.000
PF: No Magnitude 1.000 0.000
PF: No Move 1.000 0.000
PF: No S Repulsion 1.000 0.001
Appendix F Partial Observability Feature Analysis
To further understand the counter-intuitive result that limited visibility agents outperform full visibility agents (Figure 5), we conducted a feature importance analysis measuring how strongly each trained model responds to observed human positions.
Methodology. We computed Spearman correlations ($\rho$) between the mean relative positions of visible humans and the agent’s movement actions. Specifically, we measured correlations between the mean x-displacement of observed humans and the agent’s x-movement ($\rho_{\Delta x}$), and similarly for the y-dimension ($\rho_{\Delta y}$).
[…]
Interpretation. The stronger correlations under limited visibility occur because full visibility dilutes attention across all humans, many of whom are distant and pose negligible infection risk. Limited visibility effectively filters the observation space to include only nearby individuals who have a larger impact on the effective transmission rate since transmission follows distance decay.
Limited visibility thus acts as an implicit attention mechanism, restricting observations to the radius where humans actually pose significant threat. Full visibility forces agents to process information about distant humans who have near-zero impact on infection risk, increasing observation dimensionality without proportional informational benefit. These results suggest that appropriately scoped observations matched to the task’s distance-dependent mechanics produce more focused and effective policies, rather than requiring more expressive architectures to handle larger state spaces.
Table 14: Spearman correlations between mean relative positions of visible humans and agent movement actions across visibility conditions. Negative values indicate avoidance behavior. All correlations significant at $p<0.001$.
Visibility $\bm{\rho_{\Delta x}}$ $\bm{\rho_{\Delta y}}$
Limited (r=10) -0.557 -0.540
Limited (r=15) -0.441 -0.464
Limited (r=20) -0.342 -0.438
Full Visibility -0.220 -0.234
Appendix G Compute Resources
@@
Appendix H Additional Results and Experiments